Please show all the steps and explain.

Prove that every amount of postage of 18 cents or more can be formed using just 4-cent and 7-cent stamps
Answer #1

P(n): Every stamp of amount 18 cents can be made using 4 and 7 cent stamps

Base case: 18 can be made as 2 7 4 18

Induction assumption: We can make stamps of 18 <nくん using just 4 and 7 cent stamps

We will show that we can also made a stamp of k+1 cents

Note that k can be made using 4 and 7 cent stamps means there exist non-negative a and b such that k4a 7b

So that k14a 7b1 4a7b2.4 7 4(a 2)7(b- 1) is a valid way of making k+1 cent stamp

Unless b=0 in which case k4a in which we can write k14a 1 4a 3.7-5.4 4(a -5)+3 7

Which is a valid way of making k+1 stamps unless a\le 4

Note that both b=0 and a\le 4 are not possible because then k4a 76 <16 contradicting k\geq 18

So that we can always make a stamp of k+1 cents

So that by induction, we can always make a stamp of n cents when 18 using 4-cent and 7-cent stamps

\blacksquare

Hope this was helpful. Please do leave a positive rating if you liked this answer. Thanks and have a good day!